OpenStudy (anonymous):

A bicyclist coasts down a 7 degree hill at a steady speed of 5.0m/s. Assuming a total mass of 75kg (bicycle plus rider), what must be the cyclist's power output to climb the same hill at the same speed?

OpenStudy (anonymous):

The work going down equals the work going up, assuming no friction or air resistance. \[W = m \cdot g \cdot \sin(\theta) \cdot \vec v\]